### v4

@@ 3.2.3 Training Data
Triplet Data (Reference Image, Target Image, Target Caption). This dataset format is directly aligned with the PCS objectives, establishing a clear relation between the reference and the personalized content. However, such large-scale triplet samples are not widely available. Several strategies have been proposed to mitigate this issue: 1) Data Augmentation. Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data [57]. 2) Synthetic Sample Generation. Methods like SuTI [52] utilize multiple TTF models to generate synthetic samples, which are then paired with original references. 3) Utilizing Recognizable SoIs. Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization [65].
Dual Data (Reference Image, Reference Caption). This dataset is essentially a simplified version of the triplet format, where the personalized content is the original image itself. Such datasets are more accessible, including collections like LAION [66] and LAION-FACE [67]. However, a notable drawback is that training tends to focus more on reconstructing the reference image rather than incorporating the text prompts. Consequently, models trained on this type of data might struggle with complex prompts that require substantial modifications or interactions with objects.
@@ 3.3 Hybrid Framework
Recently, some works have started to explore the combination of TTF and PTA methods. HyperDreamBooth [26] states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details. They first develop a PTA network, followed by a subject-driven fine-tuning. Similarly, DreamTuner [39] pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction. Then an additional fine-tuning stage is conducted for fine identity preservation. In contrast, SuTI [52] first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
@@ 4.1 Attention-based Operation
Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively [68]. In diffusion models, these operations generally involve manipulating the way a model focuses on different parts of data, often through a method known as the Query-Key-Value (QKV) scheme. While large-scale pre-training has equipped this module with strong feature extraction capabilities, there remains significant ongoing work to enhance its performance for customization tasks.
Explicit attention weight manipulation. A cluster of studies focuses on restricting the influence of the SoI token within the attention layers. For example, Mix-of-Show [25] designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity. This avoids the misalignment between the words and visual regions. DreamTuner [39] designs a self-subject-attention layer to further refine the subject identity. This attention module takes the features of the generated image as query, the concatenation of generated features as key, and reference features as value. Layout-Control [69] adjusts attention weights specifically around the layout without additional training. Cones 2 [70] also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation. VICO [71] inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
Implicit attention guidance. In addition to these explicit attention weights modification methods, many researchers [5, 22, 38, 63, 45, 72, 37] employ localization supervision in the cross-attention module. Specifically, they train cross-attention modules using coordinate-aware loss functions, forcing attention maps to align with annotated subject positions. DreamTuner [39] further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
@@ 4.2 Mask-guided Generation
Pixel-level mask. Benefitting from advanced segmentation methods like SAM [73], the SoI can be precisely isolated from the background. Based on this strategy, plenty of studies [44, 32, 38, 45, 34, 35, 37, 74, 75] choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances. Another technique [49] further adds the masked background reconstruction for better disentanglement. In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal [5, 22, 38, 63, 45, 72, 37] to adjust the attention’s concentration adaptively. Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns [47, 76, 63].
Feature-level mask. In addition to the pixel-level manipulation, masks can be extended to feature-level operations. DisenBooth [77] defines an identity-irrelevant embedding with a learnable mask. By maximizing the cosine similarity between the identity-preservation embedding and identity-irrelevant embedding, the mask will adaptively exclude the redundant information, and thus the subject appearance can be better preserved. AnyDoor [78] defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process. Face-Diffuser [79] determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model. Each model makes its own noise prediction, and the final noise output is a composite created through mask-guided concatenation.
@@ 4.3 Data Augmentation
Compositional augmentation. Some methods enhance data diversity through classical image augmentation like blending and spatial rearrangement. SVDiff [22] manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios. Such concept composition is also used in other works [63, 80, 81]. BLIP-Diffusion [57] segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset. StyleAdapter [82] chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style. PACGen [75] shows that the spatial position entangles with the identity information. Thus rescaling, center crop, and relocation are effective augmentation solutions.
Synthetic data. Generated synthetic data can provide a large amount of training resources when coupled with quality assurance mechanisms. SuTI [52] establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI. These synthetic samples then train the target PTA model. Similarly, DreamIdentity [48] leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image. StyleDrop [24] and GAL [19] implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
External sources. It is intuitive to leveraging web resources to expand training datasets. COTI [83] adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
@@ 4.4 Regularization
Auxiliary data regularization. To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, studies start to use an additional dataset composed of images with the same category of the SoI [4]. By reconstructing these images, the personalized model is required to generate diverse instances of the class while adapting to the target subject. Building on this strategy, StyleBoost [84] introduces an auxiliary style-specific data to separate content and aesthetic adaptation. Later, a dataset [40] is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
Text embedding constraints. The semantic richness of pre-trained text (e.g., subject class name) provides a powerful regularization signal for personalized generation. By strategically constraining how subject-specific representations interact with established linguistic concepts in the embedding space, these approaches can achieve better generalization ability. For example, Perfusion [21] constrains key projections toward class noun embeddings while learning value projections from subject images. Inspired by coached active learning [85, 86], which uses anchor concepts for optimization guidance, Compositional Inversion [87] employs a set of semantically related tokens as anchors to constrain the token embedding search. In addition, some works [88, 34] regularize learnable token offsets relative to pre-trained CLIP embeddings. By minimizing the offset, the final word embedding is able to achieve better text alignment. Similarly, Cones 2 [70] minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun. And [33] optimizes the learnable token towards the mean textual embedding of 691 well-known names. Domain-Agnostic [89] proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples. On the other hand, VICO [71] empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI. To leverage this discovery, an L2 loss is leveraged to reduce the difference of attention similarity logits between the SoI token and <|EOT|>.
@@ 5.1 Personalized Object Generation
An important branch in TTF methods is to optimize the learnable token embeddings. The first work started from Textual Inversion [7], which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI. One of the significant benefits of this method is its minimal storage requirement as the new tokens consumes just a few kilobytes. However, the method compresses complex visual features into a small set of parameters, which can lead to long convergence times and a potential loss in visual fidelity. Recent work aims to address these limitations through several directions. DVAR [90] improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence. For enhanced representation, P+ [18] introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters. NeTI [30] advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers. ProSpect [31] recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process. Similarly, a study by [91] shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes. Later, HiFiTuner [32] integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference. Alternative approaches, like DreamArtist [92], choose to optimize both negative and positive prompt embeddings to refine the detail preservation. In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework [93] and gradient-free evolutionary optimization [94]. In summary, recent developments following the foundational work of Textual Inversion focus on reducing training times and enhancing the visual quality of generated images.
@@
In the realm of TTF methods for PCS, there is a clear shift towards fine-tuning model weights rather than just the token embeddings. This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data [25, 95]. DreamBooth [4] proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model. Besides, a regularization dataset containing 20-30 images with the same category as SoI is adopted to overcome the overfitting problem. These two combined approaches achieve impressive performance that largely promotes the progress of the research on image personalization. However, fine-tuning the entire model for each new object causes considerable storage costs, potentially rendering widespread application. To address this, Custom Diffusion [20] focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency. Further approach, Perfusion [21], also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding. COMCAT [96] introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs. Additionally, methods like adapters [23, 24] and LoRA variants [25, 26, 27, 34, 35, 36] are increasingly utilized in personalized generation for parameter-efficient fine-tuning. It is worth noting that the token embedding fine-tuning is compatible with diffusion weight fine-tuning. Multiple methods [37, 38, 39] have started using combined weight fine-tuning.
@@
PTA framework. For practical deployment of PCS systems, fast response time is a crucial factor. PTA methods enable real-time generation (less than 10 sec/subject) by leveraging large-scale pre-training to avoid per-subject optimization during the inference phase. Re-Imagen [97] introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt. While it is not specifically tailored for object personalization, it demonstrates the feasibility of training reference-conditioned frameworks. Later, ELITE [47] specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds. Both fused features and local features serve as conditions for the denoising process. Similarly, InstantBooth [98] retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively. Additionally, UMM-Diffusion [61] designs a multi-modal encoder that produces fused features based on the reference image and text prompt. The text features and multi-modal hidden state are seen as guidance signals to predict a mixed noise. Another work, SuTI [52], adopts the same architecture as Re-Imagen. The difference lies in the training samples which are produced by a massive number of TTF models, each tuned on a particular subject set. This strategy promotes a more precise alignment with personalization at an instance level rather than the class level of Re-Imagen. Moreover, Domain-Agnostic [89] combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token. Besides, They introduce a dual-path attention module separately conditioned on the nearest token and pseudo embedding. Compared to the methods that use separate encoders to process a single modality, some works have explored the usage of pre-trained multi-modal large language models (MLLM) that can process text and image modality within a unified framework. For example, BLIP-Diffusion [57] utilizes the pre-trained BLIP2 [62] that encodes multimodal inputs including the SoI reference and a class noun. The output embedding is then concatenated with context description and serves as a condition to generate images. Further, Customization Assistant [58] and KOSMOS-G [60] replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description. Meanwhile, to meet the standard format of Stable Diffusion, a network is trained to align the dimension of the output embedding.

### v5

@@ 3.2.3 Training Data
Triplet Data (Reference Image, Target Image, Target Caption). This dataset format is directly aligned with the PCS objectives, establishing a clear relation between the reference and the personalized content. However, such large-scale triplet samples are not widely available. Several strategies have been proposed to mitigate this issue: 1) Data Augmentation. Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data [10li2024blip]. 2) Synthetic Sample Generation. Methods like SuTI [11chen2024subject] utilize multiple TTF models to generate synthetic samples, which are then paired with original references. 3) Utilizing Recognizable SoIs. Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization [60yuan2023inserting].
Dual Data (Reference Image, Reference Caption). This dataset is essentially a simplified version of the triplet format, where the personalized content is the original image itself. Such datasets are more accessible, including collections like LAION [schuhmann2022laion] and LAION-FACE [zheng2022general]. However, a notable drawback is that training tends to focus more on reconstructing the reference image rather than incorporating the text prompts. Consequently, models trained on this type of data might struggle with complex prompts that require substantial modifications or interactions with objects.
@@ 3.3 Hybrid Framework
Recently, some works have started to explore the combination of TTF and PTA methods. HyperDreamBooth [58ruiz2023hyperdreambooth] states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details. They first develop a PTA network, followed by a subject-driven fine-tuning. Similarly, DreamTuner [86hua2023dreamtuner] pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction. Then an additional fine-tuning stage is conducted for fine identity preservation. In contrast, SuTI [11chen2024subject] first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
@@ 4.1 Attention-based Operation
Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively [hassanin2024visual]. In diffusion models, these operations generally involve manipulating the way a model focuses on different parts of data, often through a method known as the Query-Key-Value (QKV) scheme. While large-scale pre-training has equipped this module with strong feature extraction capabilities, there remains significant ongoing work to enhance its performance for customization tasks.
Explicit attention weight manipulation. A cluster of studies focuses on restricting the influence of the SoI token within the attention layers. For example, Mix-of-Show [46gu2024mix] designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity. This avoids the misalignment between the words and visual regions. DreamTuner [86hua2023dreamtuner] designs a self-subject-attention layer to further refine the subject identity. This attention module takes the features of the generated image as query, the concatenation of generated features as key, and reference features as value. Layout-Control [152chen2024training] adjusts attention weights specifically around the layout without additional training. Cones 2 [62liu2023cones] also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation. VICO [20hao2023vico] inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
Implicit attention guidance. In addition to these explicit attention weights modification methods, many researchers [54xiao2023fastcomposer, 8han2023svdiff, 74avrahami2023break, 16ma2023subject, 81safaee2023clic, 44rahman2024visual, 90hyung2023magicapture] employ localization supervision in the cross-attention module. Specifically, they train cross-attention modules using coordinate-aware loss functions, forcing attention maps to align with annotated subject positions. DreamTuner [86hua2023dreamtuner] further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
@@ 4.2 Mask-guided Generation
Pixel-level mask. Benefitting from advanced segmentation methods like SAM [kirillov2023segment], the SoI can be precisely isolated from the background. Based on this strategy, plenty of studies [21zhang2023personalize, 48wang2023hifi, 74avrahami2023break, 81safaee2023clic, 51lu2024object, 95achlioptas2023stellar, 90hyung2023magicapture, 92tang2023retrieving, 34li2023generate] choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances. Another technique [85cai2023decoupled] further adds the masked background reconstruction for better disentanglement. In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal [54xiao2023fastcomposer, 8han2023svdiff, 74avrahami2023break, 16ma2023subject, 81safaee2023clic, 44rahman2024visual, 90hyung2023magicapture] to adjust the attention’s concentration adaptively. Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns [19wei2023elite, 24jia2023taming, 16ma2023subject].
Feature-level mask. In addition to the pixel-level manipulation, masks can be extended to feature-level operations. DisenBooth [18chen2023disenbooth] defines an identity-irrelevant embedding with a learnable mask. By maximizing the cosine similarity between the identity-preservation embedding and identity-irrelevant embedding, the mask will adaptively exclude the redundant information, and thus the subject appearance can be better preserved. AnyDoor [14chen2023anydoor] defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process. Face-Diffuser [52wang2023high] determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model. Each model makes its own noise prediction, and the final noise output is a composite created through mask-guided concatenation.
@@ 4.3 Data Augmentation
Compositional augmentation. Some methods enhance data diversity through classical image augmentation like blending and spatial rearrangement. SVDiff [8han2023svdiff] manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios. Such concept composition is also used in other works [16ma2023subject, 138chen2023videodreamer, 145wang2024customvideo]. BLIP-Diffusion [10li2024blip] segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset. StyleAdapter [26wang2023styleadapter] chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style. PACGen [34li2023generate] shows that the spatial position entangles with the identity information. Thus rescaling, center crop, and relocation are effective augmentation solutions.
Synthetic data. Generated synthetic data can provide a large amount of training resources when coupled with quality assurance mechanisms. SuTI [11chen2024subject] establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI. These synthetic samples then train the target PTA model. Similarly, DreamIdentity [89chen2023dreamidentity] leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image. StyleDrop [5sohn2024styledrop] and GAL [151zhang2024generative] implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
External sources. It is intuitive to leveraging web resources to expand training datasets. COTI [3yang2023controllable] adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
@@ 4.4 Regularization
Auxiliary data regularization. To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, studies start to use an additional dataset composed of images with the same category of the SoI [2ruiz2023dreambooth]. By reconstructing these images, the personalized model is required to generate diverse instances of the class while adapting to the target subject. Building on this strategy, StyleBoost [53park2023styleboost] introduces an auxiliary style-specific data to separate content and aesthetic adaptation. Later, a dataset [57he2023data] is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
Text embedding constraints. The semantic richness of pre-trained text (e.g., subject class name) provides a powerful regularization signal for personalized generation. By strategically constraining how subject-specific representations interact with established linguistic concepts in the embedding space, these approaches can achieve better generalization ability. For example, Perfusion [7tewel2023key] constrains key projections toward class noun embeddings while learning value projections from subject images. Inspired by coached active learning [wei2012coaching, wei2011coached], which uses anchor concepts for optimization guidance, Compositional Inversion [9zhang2023compositional] employs a set of semantically related tokens as anchors to constrain the token embedding search. In addition, some works [12gal2023encoder, 51lu2024object] regularize learnable token offsets relative to pre-trained CLIP embeddings. By minimizing the offset, the final word embedding is able to achieve better text alignment. Similarly, Cones 2 [62liu2023cones] minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun. And [66pang2023cross] optimizes the learnable token towards the mean textual embedding of 691 well-known names. Domain-Agnostic [17arar2023domain] proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples. On the other hand, VICO [20hao2023vico] empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI. To leverage this discovery, an L2 loss is leveraged to reduce the difference of attention similarity logits between the SoI token and <|EOT|>.
@@ 5.1 Personalized Object Generation
An important branch in TTF methods is to optimize the learnable token embeddings. The first work started from Textual Inversion [1gal2022image], which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI. One of the significant benefits of this method is its minimal storage requirement as the new tokens consumes just a few kilobytes. However, the method compresses complex visual features into a small set of parameters, which can lead to long convergence times and a potential loss in visual fidelity. Recent work aims to address these limitations through several directions. DVAR [69voronov2024loss] improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence. For enhanced representation, P+ [23voynov2023p+] introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters. NeTI [61alaluf2023neural] advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers. ProSpect [30zhang2023prospect] recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process. Similarly, a study by [79agarwal2023image] shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes. Later, HiFiTuner [48wang2023hifi] integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference. Alternative approaches, like DreamArtist [68dong2022dreamartist], choose to optimize both negative and positive prompt embeddings to refine the detail preservation. In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework [56chae2023instructbooth] and gradient-free evolutionary optimization [73fei2023gradient]. In summary, recent developments following the foundational work of Textual Inversion focus on reducing training times and enhancing the visual quality of generated images.
@@
In the realm of TTF methods for PCS, there is a clear shift towards fine-tuning model weights rather than just the token embeddings. This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data [46gu2024mix, 80zhao2023catversion]. DreamBooth [2ruiz2023dreambooth] proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model. Besides, a regularization dataset containing 20-30 images with the same category as SoI is adopted to overcome the overfitting problem. These two combined approaches achieve impressive performance that largely promotes the progress of the research on image personalization. However, fine-tuning the entire model for each new object causes considerable storage costs, potentially rendering widespread application. To address this, Custom Diffusion [6kumari2023multi] focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency. Further approach, Perfusion [7tewel2023key], also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding. COMCAT [75xiao2023comcat] introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs. Additionally, methods like adapters [72xiang2023closer, 5sohn2024styledrop] and LoRA variants [46gu2024mix, 58ruiz2023hyperdreambooth, 115kong2024omg, 51lu2024object, 95achlioptas2023stellar, 76yeh2023navigating] are increasingly utilized in personalized generation for parameter-efficient fine-tuning. It is worth noting that the token embedding fine-tuning is compatible with diffusion weight fine-tuning. Multiple methods [90hyung2023magicapture, 74avrahami2023break, 86hua2023dreamtuner] have started using combined weight fine-tuning.
@@
PTA framework. For practical deployment of PCS systems, fast response time is a crucial factor. PTA methods enable real-time generation (less than 10 sec/subject) by leveraging large-scale pre-training to avoid per-subject optimization during the inference phase. Re-Imagen [67chen2022re] introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt. While it is not specifically tailored for object personalization, it demonstrates the feasibility of training reference-conditioned frameworks. Later, ELITE [19wei2023elite] specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds. Both fused features and local features serve as conditions for the denoising process. Similarly, InstantBooth [22shi2023instantbooth] retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively. Additionally, UMM-Diffusion [71ma2023unified] designs a multi-modal encoder that produces fused features based on the reference image and text prompt. The text features and multi-modal hidden state are seen as guidance signals to predict a mixed noise. Another work, SuTI [11chen2024subject], adopts the same architecture as Re-Imagen. The difference lies in the training samples which are produced by a massive number of TTF models, each tuned on a particular subject set. This strategy promotes a more precise alignment with personalization at an instance level rather than the class level of Re-Imagen. Moreover, Domain-Agnostic [17arar2023domain] combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token. Besides, They introduce a dual-path attention module separately conditioned on the nearest token and pseudo embedding. Compared to the methods that use separate encoders to process a single modality, some works have explored the usage of pre-trained multi-modal large language models (MLLM) that can process text and image modality within a unified framework. For example, BLIP-Diffusion [10li2024blip] utilizes the pre-trained BLIP2 [li2023blip] that encodes multimodal inputs including the SoI reference and a class noun. The output embedding is then concatenated with context description and serves as a condition to generate images. Further, Customization Assistant [84zhou2023customization] and KOSMOS-G [77pan2023kosmos] replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description. Meanwhile, to meet the standard format of Stable Diffusion, a network is trained to align the dimension of the output embedding.